A train is travelling at a speed of 90 km h −1 . Brakes are applied so as to produce a
uniform acceleration of −0.5 m s −2 . Find how far the train will go before it is
brought to rest.
1
Expert's answer
2020-06-04T10:04:44-0400

The initial speed of train is "v_0 = 90km\/h = 25 m\/s" and the acceleration is "-0.5m\/s^2". According to the kinematic equation:


"v_0 = -at"

where "t" is the time required to stop the train. Thus:


"t = -\dfrac{v_0}{a} = -\dfrac{25}{-0.5} = 50s"

The distance, traveled by train by this time is:


"d = v_0t + \dfrac{at^2}{2} = 25\cdot 50 - \dfrac{0.5\cdot 50^2}{2} = 625 m"

Answer. d = 625 m.